### v1

# Corrections to Kerr-Newman from Noncommutative Einstein-Maxwell equation

Abstract
In this letter we introduce the noncommutative geometry into the standard Einstein-Hilbert-Maxwell action via the $\partial_{t}\wedge\partial_{\phi}$ Drinfeld twist and solve the equation of motion pertubatively in the expansion of the noncommutative parameter $a$. The equation of motion, the NC Einstein-Maxwell equation, turns out to be effectively a problem in nonlinear electrodynamics where the energy-momentum tensor $T_{\mu\nu}$ obtains correction terms with three Faraday tensors $F_{\mu\nu}$. A solution with nonzero $a^{1}$ terms turns out to be the Kerr-Newman black hole modified with nonzero $g_{t\theta},g_{r\phi},g_{tr}$ and $g_{\phi\theta}$ components proportional to $a$, while the electromagnetic potential is the Seiberg-Witten expanded Kerr-Newman potential which introduces a nonzero $A_{\theta}$ term proportional to $a$.
preprint: RBI-ThPhys-2025-06
I Introduction
Noncommutative (NC) geometry is a candidate framework for probing the effects of quantum spacetime in physics. It has been studied in many papers, concerning various topics like black hole entropy [1, 2, 3], gravitational waves [4, 5], gauge theory [6, 7, 8], vacuum energy [9, 10], etc.. In this letter, we will use the Hopf algebra formalism, as outlined in [11, 2, 4] and many other references, to deform the Einstein-Hilbert-Maxwell (EHM) action and solve its equations of motion under the assumption that the solution fields have two Killing vectors, $\partial_{t}$ and $\partial_{\phi}$. We conclude that there exist infinitely many deformations of the EHM action, but all of them share a static solution which we found. In short, the (commutative) geometry of a (commutative) manifold $\mathcal{M}$ is encoded in its Hopf Algebra of vector fields $(\mathcal{U}(\Xi),m,\Delta,\epsilon,S)$, where
  • $\mathcal{U}(\Xi)$ is the universal enveloping algebra of the vector field Lie algebra $\Xi$. It is a unital associative algebra.
  • $m:\mathcal{U}(\Xi)\otimes\mathcal{U}(\Xi)\rightarrow\mathcal{U}(\Xi)$ is the canonical product of the $\mathcal{U}(\Xi)$ Hopf algebra, given as the composition of vector fields.
  • $\Delta:\mathcal{U}(\Xi)\rightarrow\mathcal{U}(\Xi)\otimes\mathcal{U}(\Xi)$ is the canonical coproduct defined as $\Delta(v)=v\otimes 1+1\otimes v$, for all $v\in\mathcal{U}(\Xi)$.
  • $\epsilon:\mathcal{U}(\Xi)\rightarrow\mathbb{C}$ is the canonical counit defined as $\epsilon(v)=0$, for all $v\in\mathcal{U}(\Xi)$.
  • $S:\mathcal{U}(\Xi)\rightarrow\mathcal{U}(\Xi)$ is the canonical $\mathcal{U}(\Xi)$ anti-automorphism defined as $S(v)=-v$, for all $v\in\mathcal{U}(\Xi)$.
Together those 5 objects satisfy various structural identities which define an abstract Hopf algebra. It is known that using a Drinfeld element 111which satisfies the technical cocycle and normalization conditions $\mathcal{F}\in\mathcal{U}(\Xi)\otimes\mathcal{U}(\Xi)$ it is possible to systematically deform the starting Hopf algebra into a new, deformed, one $(\mathcal{U}(\Xi),m,\Delta_{\mathcal{F}},\epsilon,S_{\mathcal{F}})$ with
[…]
and this deformed Hopf algebra is said to encode the noncommutative geometry of some noncommutative manifold $\mathcal{M}_{\star}$. Furthermore, it is also well known that the algebra $(A,\star)$ modules, i.e., the algebras on which the Hopf algebra acts  [^2] All of the relevant geometric modules on which the vector field Hopf algebra naturally acts can be seen as algebras - e.g., the pointwise algebra of functions on the manifold $C^{\infty}(\mathcal{M},\mathbb{C})$, the tensor algebra $\mathcal{T}$, etc. are all algebras., of the deformed Hopf algebra can be obtained from the underlying undeformed Hopf algebra’s algebra modules $(A,\cdot)$ simply by setting  [^3] Where $a_{1}\cdot a_{2}$ can be understood as a tensor product map $\cdot:A\otimes A\rightarrow A$ with $\cdot(a_{1}\otimes a_{2})=a_{1}\cdot a_{2}$.
$a_{1}\star a_{2}=\cdot\circ\mathcal{F}^{-1}\left(a_{1}\otimes a_{2}\right)\;,\quad\forall a_{1},a_{2}\in A\;.$ (2)
With the definition (2) it is easy to see that $(A,\star)$ is an algebra module of the deformed Hopf algebra because
$v\vartriangleright(a_{1}\star a_{2})=\star\left[\Delta_{\mathcal{F}}(v)\vartriangleright\left(a_{1}\otimes a_{2}\right)\right]$ (3)
holds for all $v\in\mathcal{U}(\Xi)$ and all $a_{1},a_{2}\in A$. We will use the the construction (2) and its consequence (3) to construct an action functional that is covariant under the deformed Hopf algebra obtained by deforming the $\mathcal{U}(\Xi)$ Hopf algebra with the $\partial_{t}\wedge\partial_{\phi}$ Drinfeld twist
$\mathcal{F}=e^{-i\frac{a}{2}\left(\partial_{t}\otimes\partial_{\phi}-\partial_{\phi}\otimes\partial_{t}\right)}=e^{-i\frac{a}{2}\partial_{t}\wedge\partial_{\phi}}\;.$ (4)
The Drinfeld twist (4) is of the Moyal type (and is also Abelian) which has the general form
$\mathcal{F}=e^{-\frac{i}{2}\theta^{\mu\nu}\partial_{\mu}\otimes\partial_{\nu}}\;,$ (5)
with the tensor $\theta$ being constant and antisymmetric.
II Noncommutative Einstein-Hilbert-Maxwell action
In this subsection we will obtain the noncommutative Einstein-Hilbert-Maxwell action. Let us consider the deformed Hopf algebra $(\mathcal{U}(\Xi),m,\Delta_{\mathcal{F}},\epsilon,S_{\mathcal{F}})$ of vector fields over spacetime $\mathcal{M}$ and its deformed algebra module of spacetime functions $C^{\infty}(\mathcal{M},\mathbb{C})$, defined using the Drinfeld twist (4) as above. An example action in this framework is, e.g., the NC charged scalar field and electromagnetism action on the fixed Reissner–Nordström (RN) background defined in [6]:
$\begin{split}S=\int d^{4}x\sqrt{-g}&\left[g^{\mu\nu}\cdot\hat{D}_{\mu}\hat{\phi}^{\dagger}\star\hat{D}_{\nu}\hat{\phi}-\mu^{2}\hat{\phi}^{\dagger}\star\hat{\phi}\right.\\ &\left.-\frac{1}{4}g^{\mu\nu}\cdot g^{\rho\sigma}\cdot\hat{F}_{\mu\rho}\star\hat{F}_{\nu\sigma}\right]\;,\end{split}$ (6)
with the deformed gauge covariant derivative defined as
$\hat{D}_{\mu}\hat{\phi}=\partial_{\mu}\hat{\phi}-iq\hat{A}_{\mu}\star\hat{\phi}\;,$ (7)
and with the star product
$\begin{split}\left(f\star g\right)(x)=&\cdot\left[\mathcal{F}^{-1}\left(f\otimes g\right)\right]\\ =&\cdot\left[e^{i\frac{a}{2}\partial_{t}\wedge\partial_{\phi}}\left(f\otimes g\right)\right]\\ =&f(x)\cdot g(x)+i\frac{a}{2}\left[\partial_{t}f(x)\partial_{\phi}g(x)\right.\\ &\left.-\partial_{\phi}f(x)\partial_{t}g(x)\right]+\mathcal{O}(a^{2})\;,\end{split}$ (8)
which is a direct application of (2) to the Drinfeld twist (4). It is important to comment, because it will be key to including the Einstein-Hilbert Lagrangian into the action (6), that in (6) the $\star$ product between metric tensors was replaced with a standard commutative product. That is because $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors of the RN background and as such any function of the metric obeys
$F(g_{\mu\nu})\star G(x)=F(g_{\mu\nu})\cdot G(x)$ (9)
due to (8). Additionally, since the fields in (6) are not elements of the algebra $(C^{\infty}(\mathcal{M},\mathbb{C}),\cdot)$, but of the deformed algebra $(C^{\infty}(\mathcal{M},\mathbb{C}),\star)$, the usual commutative $U(1)$ gauge symmetry is not a symmetry of (6). The only type of gauge transformations that the action (6) can be invariant to is the noncommutative $U(1)$ gauge transformation, denoted as $U(1)_{\star}$, whose infinitesimal expansion is given as
$\begin{split}\delta_{\hat{\lambda}}\hat{\phi}&=\hat{\lambda}\star\hat{\phi}\;,\\ \delta_{\hat{\lambda}}\hat{A}_{\mu}&=\partial_{\mu}\hat{\lambda}+i\left(\hat{\lambda}\star\hat{A}_{\mu}-\hat{A}_{\mu}\star\hat{\lambda}\right)\;,\end{split}$ (10)
with $\hat{\lambda}(x)$ being the NC gauged parameter. To distinguish the fields obeying NC gauge transformations with the fields obeying commutative gauge transformations, it is customary to denote NC gauge theory fields with a hat, e.g., $\hat{A}_{\mu}$, and the fields obeying usual $U(1)$ transfromation laws without the hat, e.g., $A_{\mu}$. In fact, the action (6) is symmetric to $U(1)_{\star}$ gauge transformations.
One can see that in the commutative limit $a\rightarrow 0$ the theory given by the action (6) becomes the usual electromagnetism + charged scalar field theory in the RN background and the $U(1)_{\star}$ gauge symmetry of the action reverts back to $U(1)$ gauge symmetry. A useful concept in the field of noncommutative gauge theory is the Seiberg-Witten (SW) map [15] which maps a commutative gauge transformation $(\lambda,A_{\mu},\phi)$ into a NC (with respect to the Moyal twist (5)) one $(\hat{\lambda},\hat{A}_{\mu},\hat{\phi})$ using the identification
$\begin{split}\hat{\lambda}&=\lambda+\frac{q}{2}\theta^{\mu\nu}\partial_{\mu}\lambda\cdot A_{\nu}+\mathcal{O}(a^{2})\;,\\ \hat{A}_{\mu}&=A_{\mu}-\frac{q}{2}\theta^{\rho\sigma}A_{\rho}\left(\partial_{\sigma}A_{\mu}+F_{\sigma\mu}\right)+\mathcal{O}(a^{2})\;,\\ \hat{\phi}&=\phi-\frac{q}{4}\theta^{\rho\sigma}A_{\rho}\left(\partial_{\sigma}\phi+D_{\sigma}\phi\right)+\mathcal{O}(a^{2})\;,\end{split}$ (11)
where $q$ is the electric charge of the scalar field $\phi$. By using the Seiberg-Witten map (11) and expanding the star product (8) up to first order, the authors of [6] were able to express the action (6) completely in terms of commutative fields and commutative products. The purely electromagnetic sector of the expanded action equals
$\begin{split}S\supset\int d^{4}x\sqrt{-g}&\Biggl{[}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{q}{8}\theta^{\alpha\beta}\Bigl{(}F_{\alpha\beta}F_{\mu\nu}F^{\mu\nu}\\ &-4F_{\mu\alpha}F_{\nu\beta}F^{\mu\nu}\Bigr{)}\Biggr{]}+\mathcal{O}(a^{2})\;.\end{split}$ (12)
We will obtain a similar expression for the Einstein-Hilbert-Maxwell Lagrangian.
If noncommutative geometry is a valid description of gravity coupled to classical electromagnetism in the regimes where gravity is not just a background, then the action in this case needs to have the form
$S[g,\hat{A}]=\int d^{4}x\left[\mathcal{L}^{\star}_{G}(g)+\mathcal{L}^{\star}_{EM}(g,\hat{A})\right]\;,$ (13)
where $\mathcal{L}^{\star}_{G}$ and $\mathcal{L}^{\star}_{EM}$ are the NC Lagrangian densities for the gravitational and electromagnetic sector and they need to reproduce the usual Einstein-Hilbert and Maxwell Lagrangian densities in the commutative limit. The problem is that there are infinitely many such densities. For example, both $\mathcal{L}^{\star(1)}_{EM}$ and $\mathcal{L}^{\star(2)}_{EM}$
$\begin{split}\mathcal{L}^{\star(1)}_{EM}=\sqrt{-g}\star g^{\mu\nu}\star g^{\rho\sigma}\star\hat{F}_{\mu\rho}\star\hat{F}_{\nu\sigma}\;\\ \mathcal{L}^{\star(2)}_{EM}=\sqrt{-g}\star g^{\mu\nu}\star\hat{F}_{\mu\rho}\star g^{\rho\sigma}\star\hat{F}_{\nu\sigma}\;\end{split}$ (14)
reproduce the Maxwell Lagrangian in the commutative limit. So far, the author is not aware of any criterion to prefer any one NC Lagrangian density over all other candidates. A similar problem is present in the NC gauge theory of gravity  [^4] Paper with T. Jurić and N. Kumara Athithamoole to appear on ArXiV soon where infinitely many noncommutative generalizations of geometric quantities, e.g, the NC metric, curvature scalars etc., have the correct commutative limits, but it is impossible to prefer any one choice over the others. But, it is still possible to find a solution to the equations of motion of the action (13). Namely, since all of the NC corrections to the action (13) come from the star product (8) and from the SW map (11), a solution $(g,A)$ of the expanded action whose fields have Killing vectors $\partial_{t}$ and $\partial_{\phi}$, has to be the same as the solution of the action
$S^{\text{NC}}_{\text{EHM}}[g,A]=\int d^{4}x\sqrt{-g}\left[R-\frac{1}{4}g^{\mu\nu}g^{\rho\sigma}\hat{F}_{\mu\rho}\cdot\hat{F}_{\nu\sigma}\right]$ (15)
$=\int d^{4}x\sqrt{-g}\biggl{[}R-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{q}{2}\theta^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}F^{\mu\nu}\biggr{]}+\mathcal{O}(a^{2})\;.$
with the same kind of symmetry in $g$ and $A$. This is due to the property (9) applied to $A$ and $g$. The action (15), which we refer to as the NC Einstein-Hilbert-Maxwell (NC EHM) action, can be thought of as the restriction of the action functional (13) to the part of the classical configuration space where the fields have Killing vectors $\partial_{t}$ and $\partial_{\phi}$. Additionally, the electromagnetic part of the action (15) expands to (12) (with $F_{\alpha\beta}$ terms vanishing due to the time and axial symmetry), so (15) is a nonlinear electrodynamics theory whose nonlinearity is induced by spacetime noncommutativity/quantumness. In conclusion, in a special symmetrical regime it is possible to actually determine the form of the NC action (13). Now it is just a question whether a solution with $\partial_{t}$ and $\partial_{\phi}$ Killing vectors exists for NC EHM action (15). The answer is that it does and we will find it in the following sections.
III The NC Einstein-Maxwell equation
Having obtained the noncommutative Einstein-Hilbert-Maxwell action, we can vary it and obtain the equations of motion for $g_{\mu\nu}$ and $A_{\mu}$, keeping in mind that any solution that we find has to obey the condition that $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors for the metric and for the potential. The equation of motion for the metric, up to $a^{1}$ order, is the NC Einstein equation
$G_{\mu\nu}=2T_{\mu\nu}\;,$ (16)
with
$\begin{split}&T_{\mu\nu}=\frac{1}{4}g_{\mu\nu}F_{\mu\nu}F^{\mu\nu}+F_{\mu}\,^{\rho}F_{\nu\rho}-\\ &\frac{q}{2}\theta^{\alpha\beta}\left(F_{\mu\alpha}F_{\sigma\beta}F_{\nu}\,^{\sigma}+F_{\nu\alpha}F_{\sigma\beta}F_{\mu}\,^{\sigma}\right)+\mathcal{O}(a^{2})\;.\end{split}$ (17)
On the other hand, the NC Maxwell equation is given in [6], but in our case the metric is dynamical instead of the Reissner Nordstrom background:
$\begin{split}&\partial_{\mu}F^{\mu\lambda}+\Gamma^{\rho}_{\mu\rho}F^{\mu\lambda}+q\theta^{\alpha\beta}\biggl{(}-\frac{1}{2}\partial_{\mu}\left(F_{\alpha\beta}F^{\mu\lambda}\right)\\ &-\frac{1}{2}\Gamma^{\rho}_{\mu\rho}F_{\alpha\beta}F^{\mu\lambda}+\partial_{\mu}\left(F_{\alpha}\,^{\mu}F_{\beta}\;^{\lambda}\right)-\\ &\partial_{\alpha}\left(F_{\beta_{\mu}}F^{\mu\lambda}\right)\biggr{)}+q\theta^{\alpha\lambda}\biggl{(}\frac{1}{2}\partial_{\mu}\left(F_{\alpha\nu}F^{\mu\nu}\right)+\\ &\frac{1}{2}\Gamma^{\rho}_{\mu\rho}F_{\alpha\nu}F^{\mu\nu}-\frac{1}{4}\partial_{\alpha}\left(F_{\mu\nu}F^{\mu\nu}\right)\biggr{)}+\mathcal{O}(a^{2})=0^{\lambda}\;.\end{split}$ (18)
We will approach solving (16) and (18) perturbatively in an expansion in $a$, disregarding $a^{2}$ effects. From the form of the energy-momentum tensor (17) and of the Maxwell equation (18), we can see that the commutative solutions to the Einstein-Maxwell problem do not solve the NC problem  [^5] Unless some accidental cancellation of seemingly unrelated terms in (16) and (18) happens which would keep the equations of motion robust to noncommutative corrections. That is not the case for Kerr-Newmann perturbations and there remain nonzero $a^{1}$ terms which are not solved with a purely commutative solution. in case the electromagnetic potential $A_{\mu}$ has $t$ and $\phi$ components, i.e., the electrovacua which describe spaces with both nonzero magnetic and electric fields are described with modified equations of motion. Having said that, we will search for a solution $(g,A)$ which is a perturbation of the Kerr-Newmann (KN) black hole (without assuming slow rotation)
$\begin{split}g_{\mu\nu}&=g_{\mu\nu}^{\text{KN}}+ah_{\mu\nu}\\ A_{\mu}&=A_{\mu}^{\text{KN}}+aB_{\mu}\;,\end{split}$ (19)
since the Kerr-Newmann commutative solution is such that the Einstein and Maxwell equations will have nonzero $a^{1}$ terms - which we can try to find a nontrivial solution for in terms of $h_{\mu\nu}$ and $B_{\mu}$. Plugging in the Kerr-Newmann solution
$\begin{gathered}g^{\text{KN}}_{\mu\nu}=\left[\begin{matrix}\frac{\frac{J^{2}\sin^{2}{\left(\theta\right)}}{M^{2}}-\Delta{\left(r\right)}}{\Sigma{\left(r,\theta\right)}}&0&0&-\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\\ 0&\frac{\Sigma{\left(r,\theta\right)}}{\Delta{\left(r\right)}}&0&0\\ 0&0&\Sigma{\left(r,\theta\right)}&0\\ -\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}&0&0&\frac{\left(-\frac{J^{2}\Delta{\left(r\right)}\sin^{2}{\left(\theta\right)}}{M^{2}}+\left(\frac{J^{2}}{M^{2}}+r^{2}\right)^{2}\right)\sin^{2}{\left(\theta\right)}}{\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\\ A^{\text{KN}}_{\mu}=\left[\begin{matrix}\frac{Qr}{\Sigma{\left(r,\theta\right)}}\\ 0\\ 0\\ -\frac{JQr\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\\ \end{gathered}$ (20)
with
$\begin{split}\Sigma(r,\theta)&=\frac{J^{2}\cos^{2}{\left(\theta\right)}}{M^{2}}+r^{2}\\ \Delta(r,\theta)&=\frac{J^{2}}{M^{2}}-2Mr+Q^{2}+r^{2}\;,\end{split}$ (21)
into the NC Einstein and NC Maxwell equations we find the $a^{1}$ source terms for the $h_{\mu\nu}$ and $B_{\mu}$ as the evaluation of the NC equations (16) and (18) on the commutative Kerr-Newmann solution, i.e., as the part of the equations which the KN electrovacuum (20) does not solve. In the NC Maxwell equation (18), $\lambda=r,\phi,\theta$ equations are not solved by the KN commutative ansatz and there are around 50 terms, given as rational functions, in each nonzero component. For the NC Einstein equation we find that there are only 4 simple distinct source term components arising from plugging in the Kerr-Newmann solution into (16):
$G_{\mu\nu}(g^{\text{KN}})-2T_{\mu\nu}(g^{\text{KN}})=aS_{\mu\nu}=\left[\begin{matrix}0&aS_{tr}&aS_{t\theta}&0\\ aS_{rt}&0&0&aS_{r\phi}\\ aS_{\theta t}&0&0&aS_{\theta\phi}\\ 0&aS_{\phi r}&aS_{\phi\theta}&0\end{matrix}\right]$ (22)
with
$\begin{gathered}S_{tr}=S_{rt}=-\frac{4JQ^{3}r^{2}\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)\sin{\left(2\theta\right)}\frac{\partial}{\partial\theta}\Sigma{\left(r,\theta\right)}}{M\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{r\phi}=S_{\phi r}=\frac{8J^{2}Q^{3}r^{2}\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)\left(-2\Sigma{\left(r,\theta\right)}\cos{\left(\theta\right)}+\sin{\left(\theta\right)}\frac{\partial}{\partial\theta}\Sigma{\left(r,\theta\right)}\right)\sin^{2}{\left(\theta\right)}\cos{\left(\theta\right)}}{M^{2}\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{t\theta}=S_{\theta t}=\frac{4JQ^{3}r\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)^{2}\Delta{\left(r\right)}\sin{\left(2\theta\right)}}{M\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{\phi\theta}=S_{\theta\phi}=-\frac{8J^{2}Q^{3}r\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)^{2}\Delta{\left(r\right)}\sin^{3}{\left(\theta\right)}\cos{\left(\theta\right)}}{M^{2}\Sigma^{6}{\left(r,\theta\right)}}\;.\end{gathered}$ (23)
IV Solving the NC Einstein-Maxwell equation
By inspecting the form of (22) one is inclined to try the ansatz for the noncommutative part of the metric with 4 nonzero components
$h_{\mu\nu}=\left[\begin{matrix}0&h_{tr}(r,\theta)&h_{t\theta}(r,\theta)&0\\ h_{tr}(r,\theta)&0&0&h_{r\phi}(r,\theta)\\ h_{t\theta}(r,\theta)&0&0&h_{\theta\phi}(r,\theta)\\ 0&h_{r\phi}(r,\theta)&h_{\theta\phi}(r,\theta)&0\end{matrix}\right]\;,$ (24)
but simply using (24) as the $a^{1}$ part of the ansatz (19) and plugging it in (16) will give a system of 4 partial differential equations with each equation being hundreds of terms long. So solving the system of equations in general is computationally intractible  [^6] Calculating the system of equations for the ansatz (24) took two hours on a modern PC using the open source Sympy library. unless one specifies the ansatz in much more detail. The way the author has found the solution was by using physical arguments, not by mechanical computation.
Firstly, the electromagnetic part of the solution should be the Seiberg-Witten map of the standard Kerr-Newman potential
$A_{\mu}=\left[\begin{matrix}\frac{Qr}{\Sigma{\left(r,\theta\right)}}\\ 0\\ 0\\ -\frac{JQr\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\end{matrix}\right]+a\left[\begin{matrix}0\\ 0\\ -\frac{qJQ^{2}r^{2}\sin{\left(2\theta\right)}}{2M\Sigma^{2}{\left(r,\theta\right)}}\\ 0\end{matrix}\right]\;,$ (25)
which is not surprising since the theory that we are solving is a $U(1)_{\star}$ noncommutative gauge theory. The metric part of the solution can be obtained by considering the scalar field action (6) in the (commutative) KN background, and in the manner of [7], the author searched for an effective metric $\hat{g}$ for which the Klein-Gordon operator equals the NC scalar field action in the fixed KN background
$\frac{\delta S}{\delta\phi}=\square_{\hat{g}}\phi\;.$ (26)
To find the solution $\hat{g}$ of the equation (26), the author used the metric part of the ansatz (19)
$\hat{g}_{\mu\nu}=g^{\text{KN}}_{\mu\nu}+a\hat{h}_{\mu\nu}\;,$ (27)
and by comparing second-derivative terms of $\phi$ in the right-hand and the left-hand side of (26), the unknown metric components were found to be the variables of an algebraic system of equation whose solutions are
$\begin{split}\hat{h}_{tr}&=-\frac{qJMQ\left(J^{2}\cos^{2}{\left(\theta\right)}-M^{2}r^{2}\right)\sin^{2}{\left(\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{t\theta}&=-\frac{qJM^{3}Qr\Delta{\left(r\right)}\sin{\left(2\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{r\phi}&=\frac{qQ\left(J^{2}+M^{2}r^{2}\right)\left(J^{2}\cos^{2}{\left(\theta\right)}-M^{2}r^{2}\right)\sin^{2}{\left(\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{\theta\phi}&=\frac{qJ^{2}M^{2}Qr\Delta{\left(r\right)}\sin^{2}{\left(\theta\right)}\sin{\left(2\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;.\\ \end{split}$ (28)
Incredibly, the same metric components (28), up to a factor of $-2$
$h_{\mu\nu}=-2\hat{h}_{\mu\nu}\;,$ (29)
together with the Seiberg-Witten expanded potential (25), solve the noncommutative equations (16) and (18). In conclusion, a charged rotating solution of the NC Einstein-Hilbert-Maxwell action (15) is given as
$\begin{gathered}g_{\mu\nu}=\left[\begin{matrix}\frac{\frac{J^{2}\sin^{2}{\left(\theta\right)}}{M^{2}}-\Delta{\left(r\right)}}{\Sigma{\left(r,\theta\right)}}&-2a\hat{h}_{tr}{\left(r,\theta\right)}&-2a\hat{h}_{t\theta}{\left(r,\theta\right)}&-\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\\ -2a\hat{h}_{tr}{\left(r,\theta\right)}&\frac{\Sigma{\left(r,\theta\right)}}{\Delta{\left(r\right)}}&0&-2a\hat{h}_{r\phi}{\left(r,\theta\right)}\\ -2a\hat{h}_{t\theta}{\left(r,\theta\right)}&0&\Sigma{\left(r,\theta\right)}&-2a\hat{h}_{\theta\phi}{\left(r,\theta\right)}\\ -\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}&-2\hat{h}_{r\phi}{\left(r,\theta\right)}&-2a\hat{h}_{\theta\phi}{\left(r,\theta\right)}&\frac{\left(-\frac{J^{2}\Delta{\left(r\right)}\sin^{2}{\left(\theta\right)}}{M^{2}}+\left(\frac{J^{2}}{M^{2}}+r^{2}\right)^{2}\right)\sin^{2}{\left(\theta\right)}}{\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\;,\\ A_{\mu}=\left[\begin{matrix}\frac{Qr}{\Sigma{\left(r,\theta\right)}}\\ 0\\ -\frac{JQ^{2}ar^{2}\sin{\left(2\theta\right)}}{2M\Sigma^{2}{\left(r,\theta\right)}}\\ -\frac{JQr\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\;,\end{gathered}$ (30)
and it satisfies the condition that $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors for $g_{\mu\nu}$ and $A_{\mu}$ so our solution (30) is valid for any action of the form (13) which reproduces the usual Einstein-Hilbert-Maxwell action in the commutative limit $a\rightarrow 0$.
It is worth spending a few words on some properties of the noncommutative solution (30). Firstly, its Komar mass and angular momentum are $M$ and $J$ respectively, so they remain unchanged in the noncommutative spacetime. Secondly, we can see that the dimensional parameter $q$ is present in both the metric and potential part of the solution. This dimensional parameter equals the charge of the scalar field in the theory (6), but since our theory (15) does not have a scalar field, we can not decide on a specific value of the new hair $q$. Notice that in the Seibeg-Witten map (11) the presence of the parameter $q$ is necessary because otherwise the map would not be dimensionally consistent. In the limit $J\rightarrow 0$ we obtain the same metric (up to the $-2$ factor) as the effective metric in [7] where it was also noticed that the effective metric solves the Einstein equation  [^7] In the limit $J\rightarrow 0$ the electromagnetic potential is simple enough that it does not introduce any $a^{1}$ terms to the Einstein equation or the Maxwell equation. up to $\mathcal{O}(a^{2})$.
V Final remarks
In this letter we have explored the inclusion of noncommutative geometry via the $\partial_{t}\wedge\partial_{\phi}$ coordinate Drinfeld twist into the standard Einstein-Hilbert-Maxwell action. We have found that there are infinitely many noncommutative actions which reproduce the EHM action in the commutative limit. That is because in the commutative geometry, there are also infinitely many equivalent EHM actions due to the commutativity of the multiplication, e.g.,
$\mathcal{L}_{1}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}=\mathcal{L}_{2}=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$ (31)
or, e.g., any convex combination of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ gives again an equivalent action. Still, we realized that under sufficiently symmetric conditions, all the NC actions share the equations of motion (16) and (18) which correspond to a nonlinear electrodynamics theory. We have solved those equation of motion and found a family of solutions, parametrized by $q$, which correspond to perturbative corrections to the commutative Kerr-Newman solution, with 4 new metric components, $g_{tr},g_{t\theta},g_{r\phi},g_{\phi\theta}$ appearing as well as one new electromagnetic potential component, $A_{\theta}$, appearing in the solutions. This letter is the first time, to author’s awareness, that the $\partial_{t}\wedge\partial_{\phi}$ Drinfeld twist defornation formalism was applied to rotating black hole physics. It is also the first time that in this formalism a correction to the black hole physics is obtained directly as a solution to the equation of motion of the NC action and not by some indirect methods. This letter serves as the proof of concept that the quantum nature of spacetime can indeed affect the black hole solutions. Finally, [11] established a result that commutative gravitational solutions with (at least) two Killing vectors are also solutions to the noncommutative Einstein equation constructed from a Killing Drinfeld twist. Here we give more depth to this result by proving that even if the vectors that define the twist are Killing for all of the commutative fields, they can still fail to solve the noncommutative Einstein-matter equation in case the matter content contains gauge fields complicated enough that the Seiberg-Witten map can introduce new nonzero source terms. Never the less, the claim from [11] can be slightly weakened to say that in such cases, the left-hand side of the Einstein equation remains undeformed, it is just the sources which can be deformed from the deformation of gauge symmetry.
VI Acknowledgment
This research was supported by the Croatian Science Foundation Project IP-2020-02-9614 ”Search for Quantum spacetime in Black Hole QNM spectrum and Gamma Ray Bursts”. The author would like to thank dr. Naveena Kumara Athithamoole for useful discussion.